# peapods: A Rust-Accelerated Monte Carlo Package for Ising Spin Systems

### v1

@@ Abstract
We present peapods [1] (github.com/PeaBrane/peapods), an open-source Python package for Monte Carlo simulation of Ising spin systems with arbitrary coupling constants on arbitrary-dimensional hypercubic lattices with periodic boundary conditions. The computational core is written in Rust and exposed to Python via PyO3 [2], combining the ergonomic interface of Python with the performance of compiled, memory-safe code. The package implements Metropolis [3] and Gibbs single-spin-flip algorithms, Swendsen–Wang [4] and Wolff [5] cluster updates, and parallel tempering [6]. Replica-level parallelism is achieved through the Rayon [7] work-stealing scheduler. We validate the implementation against the exact critical temperature of the two-dimensional Ising model [8] via finite-size scaling of the Binder cumulant [9].
@@ 1 Introduction
Monte Carlo simulations of spin systems are conceptually straightforward but computationally demanding. Near the critical temperature, critical slowing down causes single-spin-flip algorithms to decorrelate slowly, requiring cluster algorithms [4, 5] for efficient sampling. Disordered systems additionally benefit from parallel tempering [6] to escape metastable states.
@@
Rust offers an alternative: compiled performance competitive with C, strong memory safety guarantees enforced at compile time, and ergonomic tooling. The PyO3 library [2] and Maturin build tool [11] provide a mature ecosystem for writing Python extensions in Rust. In this work, we describe peapods, a package that places the entire sampling loop in Rust while exposing a clean Python interface for model construction, analysis, and visualization.
@@ 2 Model
We consider the Ising Hamiltonian on a $d$-dimensional hypercubic lattice $\Lambda$ with periodic boundary conditions:
@@
where $\sigma_{i}\in\{-1,+1\}$ are the spin variables, the sum runs over nearest-neighbor pairs, and $J_{ij}$ are the coupling constants. The couplings can be specified as an arbitrary user-provided array of shape $(\text{lattice\_shape},d)$, or via one of three built-in distributions:
@@
The package simulates $R$ replicas of the system at temperatures $T_{1}<T_{2}<\cdots<T_{R}$, enabling parallel tempering and the computation of temperature-dependent observables.
@@ 4.2 Lattice geometry
Rather than precomputing and storing a neighbor table, the Lattice struct stores only the lattice shape and precomputed row-major strides. The neighbor of flat index $i$ in dimension $d$ (forward or backward) is computed via modular arithmetic:
$i^{\prime}=i-c_{d}\cdot s_{d}+c_{d}^{\prime}\cdot s_{d},\qquad c_{d}=\lfloor i/s_{d}\rfloor\bmod L_{d},$ (4)
where $s_{d}$ is the stride, $L_{d}$ is the extent in dimension $d$, and $c_{d}^{\prime}=(c_{d}\pm 1)\bmod L_{d}$. This saves memory ($2dN$ integers per replica for a neighbor table vs. $d$ integers total) and is fast in practice due to the regularity of the access pattern.
@@ 5 Validation
We validate the implementation using the exactly solvable two-dimensional ferromagnetic Ising model, whose critical temperature is $T_{c}=2/\ln(1+\sqrt{2})\approx 2.269$ [8].
@@
Figure 1 shows $U_{L}(T)$ for $L=8,16,32,64$ computed with 50,000 sweeps of Metropolis updates combined with Swendsen–Wang cluster updates and parallel tempering. All four curves cross at $T\approx 2.27$ with $U\approx 0.61$, consistent with the known exact critical temperature and the universal Binder cumulant value of the 2D Ising universality class.
@@ 7 Usage
print(model.heat_capacity)
The package supports arbitrary-dimensional lattices (2D, 3D, etc.) and all three coupling types. Installation from PyPI is via pip install peapods, with pre-built wheels for Linux (x86_64, aarch64), macOS (Intel, Apple Silicon), and Windows (x86_64).
@@ 8 Conclusion
We have presented peapods, a Monte Carlo simulation package for Ising spin systems that combines a Python interface with a Rust computational backend. The package implements the standard toolkit of Metropolis, Gibbs, Swendsen–Wang, Wolff, and parallel tempering algorithms, parallelized across temperature replicas via Rayon. The implementation is validated against the exact 2D Ising critical point and offers significant performance improvements over Python/NumPy implementations. The package is open-source and available on PyPI.
@@
  • Arbitrary lattice geometries: supporting non-hypercubic lattices (triangular, honeycomb, etc.) via user-provided adjacency structures.
@@
  • Replica cluster moves: implementing multi-replica cluster algorithms such as the Houdayer and Jörg moves that exploit correlations between replicas to accelerate decorrelation in spin glasses.

### v2

@@ Abstract
We present peapods [1] (github.com/PeaBrane/peapods), an open-source Python package for Monte Carlo simulation of Ising spin systems with arbitrary coupling constants on periodic Bravais lattices with user-specified neighbor offsets. The computational core is written in Rust and exposed to Python via PyO3 [2], combining the ergonomic interface of Python with the performance of compiled, memory-safe code. The package implements Metropolis [3] and Gibbs single-spin-flip algorithms, Swendsen–Wang [4] and Wolff [5] cluster updates, parallel tempering [6], and three replica cluster moves for spin glasses: the Houdayer isoenergetic cluster move [7], the Jörg stochastic variant [8], and the Chayes–Machta–Redner (CMR) blue-bond algorithm [9]. Overlap statistics between replica pairs enable computation of the spin glass order parameter and Binder ratio. Replica-level parallelism is achieved through the Rayon [10] work-stealing scheduler. We validate the implementation against the exact critical temperatures of the two-dimensional Ising model on the square [11] and triangular [12, 13] lattices via finite-size scaling of the Binder cumulant [14].
@@ 1 Introduction
Monte Carlo simulations of spin systems are conceptually straightforward but computationally demanding. Near the critical temperature, critical slowing down causes single-spin-flip algorithms to decorrelate slowly, requiring cluster algorithms [4, 5] for efficient sampling. Disordered systems additionally benefit from parallel tempering [6] to escape metastable states, and from replica cluster moves such as the Houdayer isoenergetic cluster move [7] and its stochastic variants [8, 9] that exploit correlations between replicas at the same temperature to accelerate decorrelation in spin glasses.
@@
Rust offers an alternative: compiled performance competitive with C, strong memory safety guarantees enforced at compile time, and ergonomic tooling. The PyO3 library [2] and Maturin build tool [16] provide a mature ecosystem for writing Python extensions in Rust. In this work, we describe peapods, a package that places the entire sampling loop in Rust while exposing a clean Python interface for model construction, analysis, and visualization. The lattice geometry is fully general: users specify neighbor offsets as integer displacement vectors, enabling simulation on hypercubic, triangular, FCC, and other Bravais lattices.
@@ 2 Model
We consider the Ising Hamiltonian on a periodic Bravais lattice $\Lambda$ with $n$ neighbor directions:
@@
where $\sigma_{i}\in\{-1,+1\}$ are the spin variables, the sum runs over nearest-neighbor pairs, and $J_{ij}$ are the coupling constants. The lattice geometry is defined by a shape (e.g. $L\times L$) and a set of $n$ integer offset vectors specifying the forward neighbor directions; hypercubic lattices use unit vectors along each axis ($n=d$), while the triangular lattice uses offsets $\{(1,0),(0,1),(1,-1)\}$ ($n=3$, coordination $z=6$). The couplings can be specified as an arbitrary user-provided array of shape $(\text{lattice\_shape},n)$, or via one of three built-in distributions:
@@
The package simulates $R$ independent replicas, each consisting of a full parallel tempering ladder across temperatures $T_{1}<T_{2}<\cdots<T_{K}$, giving $R\times K$ total systems. Multiple replicas at the same temperature enable replica cluster moves and the computation of overlap statistics for spin glass order parameters.
@@ 3.3 Parallel tempering
3.4 Replica cluster moves
The package implements three replica cluster algorithms for spin glasses, all operating on pairs of replicas at the same temperature. The common framework is:
  1. 1.
    For each temperature, randomly pair the replicas. For each pair $(a,b)$, identify an eligible subgraph based on the overlap $\sigma_{i}^{(a)}\sigma_{i}^{(b)}$.
  2. 2.
    Build clusters on this subgraph (via BFS for Wolff-style single-cluster, or union-find for Swendsen–Wang-style global decomposition).
  3. 3.
    Exchange the spins in selected clusters between the two replicas: $\sigma_{i}^{(a)}\leftrightarrow\sigma_{i}^{(b)}$ for all $i$ in the cluster.
Houdayer ICM.
The Houdayer isoenergetic cluster move [7] restricts to the negative-overlap sites $\mathcal{N}=\{i:\sigma_{i}^{(a)}\neq\sigma_{i}^{(b)}\}$ and activates all bonds deterministically (probability 1). This exchange is isoenergetic: because every bond crossing the cluster boundary connects a site inside $\mathcal{N}$ (where the replicas disagree) to a site outside $\mathcal{N}$ (where they agree), the local field at each boundary site is unchanged by the swap. The move is always accepted with no Metropolis test, and the cluster growth is temperature-independent, which is particularly valuable at low temperatures where single-spin-flip moves freeze out.
Jörg move.
Jörg [8] proposed a refinement for three-dimensional systems where Houdayer clusters can span the lattice. The eligible subgraph is the same (negative-overlap sites), but bonds are activated stochastically with probability $p=1-e^{-4J_{ij}\sigma_{i}^{(a)}\sigma_{j}^{(a)}/T}$ on satisfied bonds ($J_{ij}\sigma_{i}^{(a)}\sigma_{j}^{(a)}>0$), breaking up system-spanning clusters into smaller pieces.
CMR move.
The Chayes–Machta–Redner (CMR) algorithm [9] generalizes the overlap cluster approach by bonding all same-sign overlap sites (not just negative-overlap), with stochastic bond activation at probability $p=1-e^{-4J_{ij}\sigma_{i}^{(a)}\sigma_{j}^{(a)}/T}$ on satisfied bonds. This provides a complementary decorrelation channel to the Houdayer and Jörg moves.
All three modes support both Wolff (single-cluster BFS) and Swendsen–Wang (global union-find) scan styles, selectable via the overlap_cluster_mode parameter.
3.5 Overlap statistics
For spin glass systems, the relevant order parameter is the overlap between replica pairs:
$q=\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}^{(a)}\sigma_{i}^{(b)}.$ (4)
The package computes $\langle q\rangle$, $\langle q^{2}\rangle$, and $\langle q^{4}\rangle$ over the measurement sweeps for each consecutive replica pair at each temperature. From these, the spin glass Binder ratio [14]
$g=1-\frac{\langle q^{4}\rangle}{3\langle q^{2}\rangle^{2}}$ (5)
can be used for finite-size scaling analysis of the spin glass transition, analogously to the ferromagnetic Binder cumulant. The crossing of $g(T)$ for different system sizes locates the critical temperature $T_{c}$ [18].
@@ 4.2 Lattice geometry
The Lattice struct supports arbitrary Bravais lattice geometries specified by a shape (extents along each axis) and a set of $n$ integer offset vectors defining the forward neighbor directions. For hypercubic lattices, these are the $d$ unit vectors; for the triangular lattice in 2D, they are $\{(1,0),(0,1),(1,-1)\}$. At construction time, the full neighbor table is precomputed: for each of the $N$ sites and each of the $n$ directions, both the forward and backward neighbor flat indices are stored, giving a table of $2nN$ entries (u32). Periodic boundary conditions are applied via rem_euclid on the lattice coordinates.
This precomputed table replaces per-call division and modular arithmetic with a single array lookup, which is faster in practice. The memory cost is modest: for example, 96 KB for a $64\times 64$ triangular lattice ($n=3$) or 6 MB for a $64^{3}$ cubic lattice ($n=3$).
@@ 5 Validation
We validate the implementation using exactly solvable two-dimensional ferromagnetic Ising models.
@@
Square lattice.
Figure 1 shows $U_{L}(T)$ for $L=8,16,32,64$ on the square lattice computed with 50,000 sweeps of Metropolis updates combined with Swendsen–Wang cluster updates and parallel tempering. All four curves cross at $T\approx 2.27$ with $U\approx 0.61$, consistent with the exact critical temperature $T_{c}=2/\ln(1+\sqrt{2})\approx 2.269$ [11] and the universal Binder cumulant value of the 2D Ising universality class.
Triangular lattice.
As a test of the general Bravais lattice geometry, we repeat the analysis on the triangular lattice with neighbor offsets $\{(1,0),(0,1),(1,-1)\}$ (coordination $z=6$). The exact critical temperature is $T_{c}=4/\ln 3\approx 3.641$ [12, 13]. For $L=8,16,32$ with 5,000 sweeps, the Binder cumulants cross at $T\approx 3.64$ with $U\approx 0.61$, confirming both the correctness of the neighbor table and the universality of the crossing value (same 2D Ising universality class).
@@ 7 Usage
# Triangular lattice ferromagnet
tri = Ising((32, 32), temperatures=np.linspace(3.0, 4.2, 32),
n_replicas=2,
neighbor_offsets=[[1, 0], [0, 1], [1, -1]])
tri.sample(5000, sweep_mode="metropolis",
cluster_update_interval=1, pt_interval=1)
print(tri.binder_cumulant)
# Spin glass with Houdayer ICM
sg = Ising((8, 8, 8), couplings="bimodal",
temperatures=np.linspace(0.8, 1.4, 24),
n_replicas=4)
sg.sample(10000, sweep_mode="metropolis",
pt_interval=1, houdayer_interval=1)
print(sg.sg_binder)
The package supports arbitrary Bravais lattice geometries via the neighbor_offsets parameter (defaulting to hypercubic when omitted) and all three coupling types. Installation from PyPI is via pip install peapods, with pre-built wheels for Linux (x86_64, aarch64), macOS (Intel, Apple Silicon), and Windows (x86_64).
@@ 8 Conclusion
We have presented peapods, a Monte Carlo simulation package for Ising spin systems that combines a Python interface with a Rust computational backend. The package implements the standard toolkit of Metropolis, Gibbs, Swendsen–Wang, Wolff, and parallel tempering algorithms, three replica cluster moves for spin glasses (Houdayer ICM, Jörg, and CMR), overlap statistics for spin glass order parameters, and support for arbitrary Bravais lattice geometries via user-specified neighbor offsets. Replica-level parallelism is achieved via Rayon. The implementation is validated against exact 2D Ising critical points on both the square and triangular lattices. The package is open-source and available on PyPI.